On March 1, 2017, a company paid an $20,700 premium on a 36-month insurance policy for coverage beginning on that date. Refer to that policy and fill in the blanks in the following table
Answer #1
Balance Sheet Income Statement
Prepaid Insurance Insurance expense
Accrual Basis Cash Basis Accrual Basis Cash Basis
Dec 31, 2017 $        26,000 $                     -   2017 $         10,000 $    36,000
Dec 31, 2018 $        14,000 $                     -   2018 $         12,000 $             -  
Dec 31, 2019 $          2,000 $                     -   2019 $         12,000 $             -  
Dec 31, 2020 $                 -   $                     -   2020 $            2,000 $             -  
Total $         36,000 $    36,000

Working

Balance Sheet Income Statement
Prepaid Insurance Insurance expense
Accrual Basis Accrual Basis
Dec 31, 2017 =36000-10000 2017 =36000*10/36
Dec 31, 2018 =26000-12000 2018 =36000*12/36
Dec 31, 2019 =14000-12000 2019 =36000*12/36
Dec 31, 2020 0 2020 =36000*2/36
Total $                  -  


In cash basis, expense is recorded when paid